1. Calculate the wavelength of a waveform given a period of 17μs.
  2. Find the input thermal noise voltage of a receiver with a bandwidth of 3.33 kHz, with an input resistance of 42 Ω, and a temperature of 29oC.
  3. Given the signal power of 2.3 μW and the noise power of 0.7 μW, calculate the signal-to-noise ratio.
  4. Three (3) AM broadcast stations are spaced at 18 kHz, beginning at 73 kHz. Each station is allowed to transmit modulating up to 6 kHz. Compute for the upper and lower sidebands of each station.
  5. Given a dc bias of 0.5mA and a bandwidth of 10 kHz, calculate for the noise current.




1
Expert's answer
2021-10-07T16:48:25-0400

1.

the wavelength:

"\lambda=vT"

where v is wave speed,

T is period

"\lambda=17v"


2) Noise voltage is given by, "V^2=4kT \int_{f1}^{f2} Rdf"

Where V= RMS input noise voltage, "f_2-f_1=band \space width=3.33KHz=3.33\times 10^3Hz"

k= Boltzmann constant"=1.38\times 10^{-23}J/K"

T=temperature in kelvin =29+273=302K

"R=Resistance=42\Omega"

Now, "V=(4kTR(f_2-f_1))^{0.5}=(4\times1.38\times 10^{-23}\times 302\times42\times 3.33\times 1000)^{0.5}=482.86\times 10^{-10}V"


3) Signal-to-noise ratio "=2.3\/0.7=23\/7=23:7"


4) Initial frequency fi = 73 kHz

Final frequency ff = (73 + 18) kHz = 91 kHz


Let Station 1, Station 2 and Station 3 be A, B and C respectively.

The fUSB of A = 73 kHz + fm (6 kHz) = 79 kHz

The fUSB of B = 79 kHz + fm (6 kHz) = 85 kHz

The fUSB of C = 85 kHz + fm (6 kHz) = 91 kHz


Where: fUSB is the Frequency of the Upper Side-bands, while fm is the modulating frequency at which each station is allowed to transmit.


Therefore, the frequency of the upper and lower side-bands of each station are given by:

  1. For A, fLSB = 73 kHz and fUSB = 79 kHz;
  2. For B, fLSB = 79 kHz and fUSB = 85 kHz;
  3. For C, fLSB = 85 kHz and fUSB = 91 kHz.


Note: fLSB means Frequency of the Lower Side-bands.


5) Root mean square value of the shot noise current in is given by the Schotty formula:


In="\sqrt{2\Iota{q}\Delta\Beta}"



Where:

"\Iota" =Dc current in Amperes

q= charge of an electron in Coulombs

"\Delta""\Beta" = the bandwidth in Hertz



Substituting


In= "\sqrt{2×0.5×10^{-3}×1.602×10^{-19}×10×10^3}"

"\sqrt{2×0.5×10^{-3}×1.602×10^{-19}×10×10^3}"



In =1.266"\u00d710^{-9}A"

=1.266nA